A stone is dropped from the top of a 45m high building. How fast will be moving when it reaches the ground?And what will its velocity be?
1
Expert's answer
2019-12-09T11:15:12-0500

height , S = 45 m


Initial velocity , U = 0

let final velocity = V


From the laws of kinematics -


"V^2 - U^2 = 2gs"

"\implies V^2 - 0 = 2\times 9.8 \times 45"

"\implies V^2 = 882"

"\implies V = 29.7 m/sec"


It reaches the ground at 29.7 m/sec